INFRARED/RAMAN (VIBRATIONAL) AND EPR SPECTRAL STUDIES OF Cu(II) HALIDE COMPLEXES WITH 3-METHYLANILINE AND 4-METHYLANILINE

\underline{M. KUMRU} and S. GUNER, Fatih University, Faculty of Arts and Sciences, Department of Physics, 34500 Büyükcekmece, Istanbul - TURKEY.